Suppose the numbers of hours of television watched by an adult in one month varies inversely as the number of books read by that adult in that month. Last month, Mr. Brady read 18 books and watched a total of 3 hours of television. This month, he read only 10 books. How many hours did he watch television?
Thank you for taking the time to answer my question.

Answer by stanbon(48558)

Hrs. TV = k/(books read)
We need to find "k".
3=k(18)
k=1/6
EQATION:
Hrs TV = (1/6)(books read)
Now to answer the question:
Hrs TV = (1/6)(10)
TV = (5/3) hrs = (5/3)(60 minutes)
TV = 100 minutes or 1 hr and 40 minutes
Cheers,
stan H.